Anyone heard about this dude? (pavelprorok on youtube)

Status
Not open for further replies.
A million dollars going begging and someone has higher priorities than to claim that million dollars. I wish I had that life.

I have heard this story before. In this forum.
 
Pavel, let's assume for a second that we're claiming the same ability.

Here's your proposal:


Propose of the protocol to test my claiming powers.

Based on the testing with 15 sets of photos (by 3 photo in each set,) 15 sets by 2 runs, in order to reach 30 sets per trial.​

Without looking in any possible- physical ways in side of the sealed envelope, just by holding it between my hands, I will be able to identify the photo that will be inside and will name it.

<snip 3,742 word description of guessing game>


Here's mine:

Proposal of protocol to test my psychic powers.​

Without looking in any possible- physical ways in side of the sealed envelope, just by holding it between my hands, I will be able to identify the photo that will be inside and will name it.

1. At a time and place convenient to the JREF and myself, a number of sealed envelopes containing photographs will be handed to me. The number of envelopes can be decided after discussion, but I would think that somewhere between 20 and 50 should suffice.

2. I will identify the photographs by holding each envelope in turn and writing down my answers.

3. To avoid ambiguous descriptions, the photographs should be of familiar, easily-described objects or landmarks. For example: A picture of a horse will be described as "a horse", otherwise it's a miss. A picture of Lenin's Tomb" will be descibed as "Lenin's Tomb". "A building" would be a miss.

4. Undoubtedly, I will make mistakes, but because I am confident in my power, I know that the results will speak for themselves, in terms of being far better than could possibly be ascribed to chance. Independent statistical anaylsis of the resullt is welcomed, but should not be required.


OK, in terms of JREF MDC standards that's a fairly pathetic protocol, and would need some considerable adjustment to be acceptable for a claim on the Megabuck, but it has one huge advantage over your "propose of the protocol to test my claiming powers." It actually matches the claim I'm making.

Mate, neither of us are statisticians or test protocol designers. We just have this amazing, easily-recognisable ability that is apparently worth a million dollars to an organisation that will actually help us prove it to the world. That's the thing about the MDC; it can never disprove supernatural abilty, only prove it, and we should use them to help us do just that.

So either give all this hair-splitting calculation of odds away and just perform as claimed or adjust the claim to fit the guessing game that you keep proposing.

Better still, I've already said how impressed I am with your ability to speak Russian and English, so why not use that real-world ability and teach us some Russian. We might even be able to help you brush up on your English. You could work away on your photo identification skills quietly in the background while taking part in other discussions here and perhaps gather a few ideas about your own abilities by reading the claims of others.

I really want to like you Pavel, for reasons unknown to me, but you're being silly.
 
Pavel is Maxwell Smart ?

Cmdr Drury: Smart, you appear to be in pretty good shape, do you work out?
Maxwell Smart: I jog one hundred miles every day! Would you believe it? One hundred miles!
Cmdr. Drury: I find that hard to believe.
Maxwell Smart: Would you believe fifty?
Cmdr. Drury: No!
Maxwell Smart: How about two push ups and a deep breath? :D
 
Cmdr Drury: Smart, you appear to be in pretty good shape, do you work out?
Maxwell Smart: I jog one hundred miles every day! Would you believe it? One hundred miles!
Cmdr. Drury: I find that hard to believe.
Maxwell Smart: Would you believe fifty?
Cmdr. Drury: No!
Maxwell Smart: How about two push ups and a deep breath? :D

Very nice of you ( and the others) to take a piss… Why don’t you just open a new thread.. something like.. “Take a piss of Pavel…” or “Lets make fun of Pavel…” or something…
 
Well, I'm convinced. This is another test that will never take place.

As long as he doesn't keep self testing, then there will be a test. If a protocol is suggested and he self tests, he will realise he can't do it. Sorry Pavel, but you started this thread with very bold, broad statements, and ever since then you have made excuses and caveats as to why you can't do what you say you can do. I do sincerely hope you get over the cognitive dissonance and realise that you really don't have powers. You could then join the JREF as a reformed believer, which there are plenty of here.

pavel_do said:
“Take a piss of Pavel…”

FYI, it's "take the piss". Just so you know. I'd like to think you're getting something out of posting on this forum.
 
As long as he doesn't keep self testing, then there will be a test. If a protocol is suggested and he self tests, he will realise he can't do it. Sorry Pavel, but you started this thread with very bold, broad statements, and ever since then you have made excuses and caveats as to why you can't do what you say you can do. I do sincerely hope you get over the cognitive dissonance and realise that you really don't have powers. You could then join the JREF as a reformed believer, which there are plenty of here.



FYI, it's "take the piss". Just so you know. I'd like to think you're getting something out of posting on this forum.

Well, thank you for the lesson. So the new thread can be open like “take the piss” =)

Re to the testing, whether you all believe in it or not.. but I do have the powers that I am claiming and I will show it during the test. I know only time will tell, but hopefully it will tell us soon. As tomorrow finale TAM is over and I can start talking to JREF and find the best way for testing my “powers”.
Have a nice weekend everybody.
 
This is a good sign Pavel. Friends take the piss (or "pull the piss) out of each other all the time, and I believe you have the sense of humour not to be offended. :)

I'd like to think, however, that we mainly take the piss out of ideas, not people, and if I've offended you on a personal level, then I apologise.

You have a nice weekend too mate. Sincerely.


Dave
 
Perhaps an interesting experiment would be for there to be 5 envelopes, two with one picture each and three with a blank sheet of the same size & weight paper used for the pictures. the envelopes would be numbered one through 5 & Pavel would have to identify what is in what envelope. Identifying a blank sheet as a photograph would obviously count as a miss. I have not calculated how many runs of this set up would be needed for 1000 to 1 odds, but it is an option based on Pavel's description of his ability. I would be curious to see how many times a envelope containing a blank sheet gets identified as a photograph.
 
Or, the other way round. He has to find the envelope with the blank in it. 5 Envelopes, 4 with different pictures in, and one with a blank sheet of paper. If pavel can identify the blank sheet 7 times out of 10 then he wins?

Surely that's easier than identifying a photo, pavel?
 
Perhaps an interesting experiment would be for there to be 5 envelopes, two with one picture each and three with a blank sheet of the same size & weight paper used for the pictures. the envelopes would be numbered one through 5 & Pavel would have to identify what is in what envelope. Identifying a blank sheet as a photograph would obviously count as a miss. I have not calculated how many runs of this set up would be needed for 1000 to 1 odds, but it is an option based on Pavel's description of his ability. I would be curious to see how many times a envelope containing a blank sheet gets identified as a photograph.

Or, the other way round. He has to find the envelope with the blank in it. 5 Envelopes, 4 with different pictures in, and one with a blank sheet of paper. If pavel can identify the blank sheet 7 times out of 10 then he wins?

Surely that's easier than identifying a photo, pavel?


How about the other way around…? What do you think? Some time a go Remiev, offered a test where there would be 9 same photos and 1 different and I had to identify the ONE =). As I have explained that would not be the best way for the reasons that I also have given. But there is an option to make it the other way… We have 10 deferent photos, (on my choice) I hold the, one by one and choose out of 10 only 2 (or 1) that I am sure the most and the vision of which was the brightest.. so I choose it and name it and we open it immediately. Then new set of 10 will be passed to me to do all the same… ( one of the envelopes even can be “empty” with the blank card) If JREF would consider it and in there opinion it could be a good, workable test. So I would work on it more. To get through the 10 envelopes does not take allot of time.
Any one can calculate the odds and how many do I need to get right out of 10 runs? ( with both cases where I would pull out 1 and 2 out of 10) And still I would like to have not 1 trial but minimum 3.
 
Hmm, I think that would only work if you specified beforehand which photo you will find, e.g. Here is set #1, find the boat picture.

If you are allowed to specify which out of the 10 you have found then it would change the odds. Can someone here work out the odds of each?
 
Hmm, I think that would only work if you specified beforehand which photo you will find, e.g. Here is set #1, find the boat picture.

If you are allowed to specify which out of the 10 you have found then it would change the odds. Can someone here work out the odds of each?

The bigetst point in this way for me, is not to be attached to a special goal ..like to find a ship.. I would like to name any 2 or 1.. that I ma sure the most. I will explain why I would not like to look for 1 specific.. ( I know every one will say.. lame excuses..) but if I would look for let say a SHIP, I would constantly repeat to my self.. Where is the ship, is this a ship etc…=) and my brain will start to draw the shop for me in my “inner vision” so when I don’t know what I am looking for.. I am more relaxed and just take what I see as I will not know the order of the photos. I don’t think that looking for 2 or 1 specific or naming any 2 or 1 out of 10, is making difference.. as odds will be the same.. 2 or 1 out of 10 in each run…

Is it possible to calculate also for 3 out of 10 per run and for 10 runs per trial, how many time do i need to get right to beat 1.000 odds.

p.s.
I will have a coffee now and will make 10 runs, by 10 photos that I mentioned so I will see my results and will see how much time it will take.. ( will post the results as soon as I am done)
 
Last edited:
I made 10 runs, with 5 minute brake between each 5. Out of each 10 I pulled out 2 that I was most sure off. Here is the results. Out of 10 runs, 4 times I got both 2 correct, 5 times only 1 out of 2 correct and 1 time both of them was wrong. The results are honest, I know they not the best but it was first time I done it 10 times. It took me in all 48 min. Seem like some of the photos in set of 10 I would need to replace, in order to remove the once that make me confused.
Just out of curiosity if it is not a hard task.. what are the odds that I beaten with the results that I given above.?
 
Pavel, when you say "I pulled out 2 that I was most sure of," do you mean

1. You guessed more than 2 of the 10 (checking to see if you were correct or not) and are reporting the results of the 2 you were most sure of, or

2. You tried to see into each envelope, but only guessed and checked to see if you were correct or not for 2 of the 10 that you felt most sure of?

I think you mean #2 above, but I just want to make sure.
 
Pavel, when you say "I pulled out 2 that I was most sure of," do you mean

1. You guessed more than 2 of the 10 (checking to see if you were correct or not) and are reporting the results of the 2 you were most sure of, or

2. You tried to see into each envelope, but only guessed and checked to see if you were correct or not for 2 of the 10 that you felt most sure of?

I think you mean #2 above, but I just want to make sure.

That is correct #2.
I "looked" in almost every envelope ( I mean held in my hands), some times not through all 10 when after a few I had a bright visions. But what ever the way.. ONLY 2 out of 10 was NAMED and OPENED to reveal the result.
 
I made 10 runs, with 5 minute brake between each 5. Out of each 10 I pulled out 2 that I was most sure off. Here is the results. Out of 10 runs, 4 times I got both 2 correct, 5 times only 1 out of 2 correct and 1 time both of them was wrong.

That sounds like an excellent result. So out of ten runs, picking 2 envelopes each time:

2/2 - 40%
1/2 - 50%
0/2 - 10%

So if we said that you can do the same thing but with slimmer odds, e.g. 5 runs and you can get both correct twice out of 5 then it should be good enough for the prelim.

I'll have a stab at the odds:

To get 2/10 on a run = 1/10 + 1/9 = 0.1 + 0.11 = 0.21
5 tries = 5 * 0.21 = 1.05

Wait, that can't be right. Damn it, can someone help me out? Are the first not mutially excusive?

To get 2/10 = 1/10 * 1/9 = 0.1 * 0.11 = 0.011
5 tries = 5 * 0.011 = 0.055

Hmm, still not enough. I think I'm getting this a bit wrong here. My A-Level stats teacher was a bit nuts.
 
That sounds like an excellent result. So out of ten runs, picking 2 envelopes each time:

2/2 - 40%
1/2 - 50%
0/2 - 10%

So if we said that you can do the same thing but with slimmer odds, e.g. 5 runs and you can get both correct twice out of 5 then it should be good enough for the prelim.

I'll have a stab at the odds:

To get 2/10 on a run = 1/10 + 1/9 = 0.1 + 0.11 = 0.21
5 tries = 5 * 0.21 = 1.05

Wait, that can't be right. Damn it, can someone help me out? Are the first not mutially excusive?

To get 2/10 = 1/10 * 1/9 = 0.1 * 0.11 = 0.011
5 tries = 5 * 0.011 = 0.055

Hmm, still not enough. I think I'm getting this a bit wrong here. My A-Level stats teacher was a bit nuts.

First, Pavel needs to say what counts as a success. I think it should be that getting any right out of ten counts as a "hit" and that X or more hits out of Y runs is a success for the MDC.

The odds of getting both wrong from the set of 10 is
(9/10)*(8/9) = 0.8 (or something close to that).

The probability of getting a hit is therefore 0.2 (roughly).

The odds of getting X>=9 hits out of Y=10 runs is about 5 in 1,000,000.
 
Well done Pavel. I can't work out odds and things but it sounds pretty impressive so far. The great thing is that this sounds much more like your original claim, and it has the advantage of a much simpler test.

The idea of just working with the envelopes you feel sure of sounds reasonable. It makes sense that some pictures could be clearer to you than others, so I hope this can be accommodated.

Cheers,

Dave.
 
Well done Pavel. I can't work out odds and things but it sounds pretty impressive so far. The great thing is that this sounds much more like your original claim, and it has the advantage of a much simpler test.

The idea of just working with the envelopes you feel sure of sounds reasonable. It makes sense that some pictures could be clearer to you than others, so I hope this can be accommodated.

Cheers,

Dave.

Thank you guys, sound like a bit of support finale..:) (touch wood :))
It seem like a not bad test for me at least.. I know for me would be better drop all the others and concentrate on one and do it for a week at least.. to sum up the results.. I don’t know how JREF would like this test.. But we shell find out soon.. as TAM is over, just don’t know whether I have to write them or they will get back to me as soon as they free… as my previous letter was considered as “bugging”
 
First, Pavel needs to say what counts as a success. I think it should be that getting any right out of ten counts as a "hit" and that X or more hits out of Y runs is a success for the MDC.

The odds of getting both wrong from the set of 10 is
(9/10)*(8/9) = 0.8 (or something close to that).

The probability of getting a hit is therefore 0.2 (roughly).

The odds of getting X>=9 hits out of Y=10 runs is about 5 in 1,000,000.

Hi Dick,

Success in what sense? As a test in general? Well as I claimed a minimum of 65%... than getting 6/7 correct out of 10 runs as cant get 6 and a half;) ) would be a success for me I guess...:) or you mean what do I consider as a “hit”? Well in this case it is obvious I need to get the photos right.. at least one out of 2 that was chosen out of 10.. but today as I said I happen to make 2 wrong “guesses” with both photos that I was sure of. To be honest I have already put those photos a side.. and will replace them with some others.. but it is bloody hard to make them all completely different.. like today 3 times I was holding the photo with the “Pirate’s flag” and sew a ship in my vision and named it as a ship and all 3 times was wrong.. as I was saying before some times I see related visions that in one or the other way, represent photo.. So I removed it now… and will change it for something elts… :) Have to be honest too, 1 time when I got 2 of them correct.. I got them in deferent order.. I don’t know but it happens to me for some reasons and some times. Still both that I named was there, just that not in order that I thought.
 
Last edited:
Status
Not open for further replies.

ISF - Join now!

Every member here is approved by hand. No bots, no spam, just people who care about evidence and honest debate.

Membership is free!

Create your free account

Back
Top Bottom